# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 1 Introduction
Numerical linear algebra (NLA) is a cornerstone of applied mathematics, scientific computing, and data science, with nascent algorithmic implementations dating back to the beginning of recorded history. Algorithms to accomplish NLA tasks, such as solving linear systems, computing eigenvalues and singular values, orthogonalizing and factorizing matrices, among many others, have been used in applications ranging from ancient Babylonian land surveying (Brezinski2022_NLA, Chapter 2) to Nobel Prize-winning work on computer vision AlexNet_2012. Throughout its long history, NLA algorithmic development has been driven by two competing imperatives, accuracy and efficiency, with the extant challenge to strike the right balance between them.
Origins of RNLA   The relatively recent emergence of randomized algorithms marks a major paradigm shift in NLA. Though Monte Carlo methods developed by Ulam and von Neumann as part of the Manhattan Project were made public in the late 1940s and 1950s Metropolis1987; Ulam1949; vonNeumann1951techniques, for decades, the NLA community was largely skeptical of randomization. Numerical analysts questioned the lack of theoretical guarantees for the accuracy of outputs of randomized algorithms, as well as their reproducibility. The general consensus was that the increase in efficiency afforded by randomization meant a sacrifice in accuracy that made the algorithms inherently unreliable, particularly because Monte Carlo methods were often inaccurate as a byproduct of the central limit theorem, cf. (martinsson2020, Section 1.2).
However, since the 1980s, that skepticism has been replaced by an abundance of work to develop randomized algorithms with provable accuracy and practical efficiency, a subfield of NLA that has grown into a research area in its own right, known as randomized numerical linear algebra (RNLA). More details on historical milestones in the RNLA methods of the 20th century can be found in halko2011; Kannan2017; Mahoney2011Randomized; martinsson2020; Tropp2015_MatrixConcentration; woodruff2015.
@@
RNLA in the 21st century   By the early 2000s, randomization was being utilized in ground-breaking applications, such as randomized sampling for web-searching AchlioptasFiatKarlinMcSherry2001; ChakrabartiDomGibsonKleinbergKumarRaghavanRajagopalanTomkins1999, including the notable Google PageRank algorithm BrinPage1998Anatomy; HenzingerHeydonMitzenmacherNajork1999; PageBrinMotwaniWinograd1998. Randomized sampling was also used to approximate matrix operations and compute low-rank matrix decompositions Achlioptas2001; Frieze2004; papadimitriou2000latent. By the mid-2000s, a unifying framework for RNLA had begun to emerge.
@@
Over the past two decades, RNLA has developed into a firmly established research area, garnering widespread attention from both practitioners and theorists. In 2006, a three-part series laid the theoretical groundwork for randomized sampling in approximate matrix multiplication DrineasKannanMahoney2006, low-rank matrix decomposition (QR and SVD) DrineasKannanMahoney2006b, and CUR decomposition DrineasKannanMahoney2006c. The same year, the separate work of DeshpandeRademacherVempalaWang2006; Deshpande2006AdaptiveSA investigated adaptive sampling for low-rank approximation, and in sarlos2006improved, a relative error SVD algorithm was proposed that utilized randomized subspace embeddings.
@@
Numerical analysts and theoretical computer scientists also provided compelling empirical evidence of the improved performance of RNLA algorithms for practical applications Deshpande2006AdaptiveSA; Friedland2005FastML; Frieze2004; Liberty2007; martinsson2006interpolation; RokhlinTygert2008. These results inspired subsequent investigations of randomized algorithms expressly for linear algebra tasks, such as preconditioning or rangefinding, paired with strong probabilistic guarantees Avron2010; HalkoMartinssonShkolniskyTygert2011; halko2011; MartinssonRokhlinTygert2011; Rokhlin2010. Twenty years later, the scope of research in RNLA now encompasses nearly every aspect of linear algebra, including eigenvalue problems bhattacharjee2023; BravermanKrishnanMusco2022; KressnerPlestenjak2024; LinSaadYang2016; Tropp2025_OSI; Tropp2022; YuanGuLi2018; general linear solvers AvronDruinskyGupta2015; Avron2010; BaboulinLiRouet2015; BalabanovGrigori2025; ChenEpperlyTroppWebber2025; NakatsukasaTropp2024; orthogonalization BalabanovGrigori2022; DuerschGu2017; FukayaKannanNakatsukasaYamamotoYanagisawa2020; GrigoriTimsit2024; JangGrigori2025; matrix function approximation AmselChenGreenbaumMuscoMusco2024; BurkeGuttel2024; ChenGreenbaumMuscoMusco2023; CortinovisKressnerNakatsukasa2024; GuttelSchweitzer2023; PerssonKressner2023; UbaruChenSaad2017; trace estimation AvronToledo2011; GrattonTitley2018; HallmanIpsenSaibaba2023; PerssonCortinovisKressner2022; RoostaKhorasaniAscher2015; UbaruSaad2018; and many, many more.
@@
Why the R in RNLA? The rise in popularity of randomized algorithms is due in large part to the commensurate progress made in computing technology. In June 2006, the performance of the world’s top supercomputer BlueGene/L was measured at 2.806 Peta-flop/s (floating point operations per second) on the High Performance Linpack (HPL) benchmark, a program that solves a large dense system of linear equations by LU with partial pivoting. As of November 2025, that rate has grown exponentially to 1.742 Exa-flop/s, achieved by today’s top supercomputer El Capitan top500list.
@@
Modern computing architectures have made it possible to solve increasingly large-scale problems; however, as evidenced by both theory and practice, the performance of classical algorithms is limited by asymptotic complexity and the “curse of dimensionality” bellman1961adaptive. As a point of reference, the amount of data collected by the Event Horizon Telescope Collaboration to generate the first image of a black hole in 2019 was over 5 Petabytes Whitwam2019, which is almost the full storage capacity of El Capitan. In other words, classical matrix algorithms with cubic cost in the input size, such as orthogonalization, may be untenable given the large-scale problems that are now considered in scientific investigations. For the best practical performance, it is necessary to develop algorithms and implementations that can fully leverage computing resources.
@@
Perhaps the greatest benefit of randomization is that it gives us the ability to restructure algorithms to attenuate computational demands on whichever computing resources most hinder performance, such as flops, matrix access, or data movement. For example, randomization admits algorithms that require just a single pass over the input data halko2011; LiYin2020SinglePassLU; Mahoney2011Randomized; Murray2023RandNLA; Tropp2017Practical; Tropp2018-MorePracticalSketching; TroppStreaming2019; Woolfe2008FastRandomized, which is especially advantageous for large-scale problems or streaming data. Randomization can also significantly reduce the number of arithmetic operations required by classical algorithms; e.g., for dense $m\times n$ overdetermined least squares problems with $n\ll m$, randomization yields solutions in roughly $O(mn+n^{3})$ operations vs. $O(mn^{2})$ with classical approaches Avron2010; Epperly2024; LacottePilanci2021; OzaslanPilanciArikan2019; PilanciWainwright2016; RokhlinTygert2008. In short, randomized algorithms produce accurate solutions with high probability while reducing working storage and communication costs, offering improved asymptotic complexity, admitting straightforward parallelization and GPU acceleration, and enabling practical investigations that would be otherwise impossible. For an expanded discussion, see (martinsson2020, Sections 1.3-1.4).
@@
RNLA for tensors   The successful incorporation of randomization into classical matrix algorithms inspired similar work to exploit randomization for fast tensor computations. Originally proposed in 1927 as a means of working with multi-way physical and chemical measurements Hitchcock1927TheEO; Hitchcock1927Multiple, tensors are now ubiquitous in scientific computing and data science. Applications that involve tensorial data include 3D image reconstruction Bendory2023Autocorrelation; Ghosh2022; Kam1980Reconstruction; Levin2018AbInitio; Zhang2024MomentMetrics; signal processing Chen2021_WirelessComm; CichockMandicLathauweri2015; Comon2001TensorDecompositions; Marot2008TensorDenoising; Miron2020; Sidiropoulos2017; quantum chemistry and physics Bischoff2012ManyBodyWaveFunctions; Khoromskaia2015TensorNM; Peng2020; Pierce2021; Szalay2015; machine learning Aidini2023FewShotTensorCompletion; Hao2013; Ji2019_ML; Panagakis2024; Tao2005; Zhou2016GroupComponentAnalysis; high-dimensional partial differential equations (PDEs) BachmayrSchneiderUschmajew2016; BoelensVenturiTartakovsky2018; DektorRodgersVenturi2021; Khoromskij2015TensorPDEs; LubichOseledetsVandereycken2015; among many others. For detailed surveys of tensors and their applications, we refer the reader to Ballard_Kolda_2025; FriedlandTammali2015LowRankApprox; Grasedyck2013; KoldaBader2009; papalexakis2017.
@@
Unfortunately, tensor practitioners suffer harsher consequences of the curse of dimensionality. For a $d$-way array of uniform size $N$, the storage and computational costs scale as $N^{d}$, quickly becoming unmanageable. As such, randomization has been essential to the development of efficient algorithms for tensor computations, popularized by several foundational works in the mid-2000s CaiafaCichocki2009; delaVega2005; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR. Two decades later, tensor practitioners increasingly rely on randomization for efficient tensor computations that are backed by the strong accuracy guarantees of well-established randomized matrix algorithms. Randomized algorithms have been developed for nearly every type of low-rank tensor decomposition in the literature, including the CP format BattaglinoBallardKolda2018; Biagioni2015; Erichson_2020; MaSolomonik2021FastAccurateRandomizedTensor; Reynolds2016; VervlietDeLathauwer2016RandomizedBlockSampling; Wang2023, the Tucker format AhmadiAsl2021Randomized; CheWeiYan2025EfficientTucker; KressnerPerisa2017RecompressionHadamard; Minster2024; MinsterSaibabaKilmer2020; Tsourakakis2010; Zhou2014, the Tensor Train format CheWei2019; CheWeiYan2024; Fonat2019; HuberSchneiderWolf2017RandomizedTT_SVD; LiYuBatselier2022FasterTensorTrainSparse, as well as other specialized representations BallaniGrasedyckKluge2013BlackBoxHTucker; BatselierYuDanielWong2018TensorNetworkRandomizedSVD; CastilloHaddockHartsockEtAl2025RandomizedKaczmarzTPRODUCT; CheWeiYan2024SketchBasedMultiplicativeUpdating; Robeva2016; TarzanaghMichailidis2018; WangCuiLi2024SVD_TensorWheel; Yuan2019; ZhangSaibabaKilmerAeron2018RandomizedTensorSVD; Zhou2014NTD.
Outline   In the following sections, we highlight recent computational advances in randomized algorithms for low-rank matrix and tensor decompositions.
Sections 1-4 focus on the matrix setting. We begin with a review of requisite linear algebra material and classical matrix decomposition algorithms. We next survey existing methods of randomized dimension reduction in addition to promising new developments in randomized sketching and sampling. We end the first part of our manuscript by discussing recent work on randomized algorithms for low-rank matrix decompositions, which are the foundation of our randomized algorithms for tensors in part two.
Sections 5-6 contain the necessary tensor prerequisites, as well as overviews of the two major tensor formats that we consider in our work: CP and Tucker. We then summarize key developments in randomized algorithms to compute low-rank tensor decompositions in these formats.
We close with some final thoughts and an outlook on the future of RNLA.
Scope   Our intent with this survey is to bring attention to important recent developments in the fast-growing body of work on RNLA. This rapid progress means that regular appraisal of the current research landscape is important for guiding new work, drawing connections to prior work, and keeping RNLA practitioners abreast of major achievements and open problems; however, it also means that important subject areas and contributions will necessarily be omitted.
@@
We quickly attempt to rectify this by pointing the interested reader to a few references on topics beyond the scope of our work. For matrices, we do not consider the vast body of work on randomized algorithms for eigenvalues and eigenvalue problems (see, e.g., HeKressnerPlestenjak2025; NakatsukasaTropp2024; Schneider2024_phd) or optimization (see, e.g., derezinski2020debiasing; na2023hessian). We also suppress the details of most theoretical results in favor of more background material to make our presentation self-contained; see, e.g., Camano2025_OSI; Kannan2017; martinsson2020. We refer the reader to Murray2023RandNLA for a survey of RNLA that centers on software implementations, and to DerezinskiMahoney2024 for a survey of RNLA from the perspective of machine learning. For tensors, we do not consider the tensor train format or other more specialized tensor decompositions and refer the reader to Ballard_Kolda_2025; Grasedyck2013.
@@ 2 Matrix Preliminaries
In this section, we summarize notation and key concepts in linear algebra that are frequently referenced in our work. We introduce our notational conventions in Sect. 2.1 and summarize several fundamental matrix decompositions and algorithms to compute them in Sect. 2.2-2.4.
@@ 2.1 Notation
To refer to coordinates of vectors, we use parentheses with subscripts, i.e. $(\bm{\mathsf{x}})_{i}$ refers to the $i$th coordinate of $\bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptn}$. We also adopt the notation of Golub and Van Loan golub2013 to reference submatrices. Namely, if $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, and $I=[i_{1},i_{2},\dots,i_{k}]\subset[m]$ and $J=[j_{1},j_{2},\dots,j_{l}]\subset[n]$ are (row and column, resp.) index sets, then $\bm{\mathsf{A}}(I,J)$ denotes the $k\times l$ matrix
$\bm{\mathsf{A}}(I,J)=\begin{bmatrix}\bm{\mathsf{A}}(i_{1},j_{1})&\bm{\mathsf{A}}(i_{1},j_{2})&\dots&\bm{\mathsf{A}}(i_{1},j_{l})\\ \bm{\mathsf{A}}(i_{2},j_{1})&\bm{\mathsf{A}}(i_{2},j_{2})&\dots&\bm{\mathsf{A}}(i_{2},j_{l})\\ \vdots&\vdots&&\vdots\\ \bm{\mathsf{A}}(i_{k},j_{1})&\bm{\mathsf{A}}(i_{k},j_{2})&\dots&\bm{\mathsf{A}}(i_{k},j_{l})\\ \end{bmatrix}$.

### v2

@@ 1 Introduction
Numerical linear algebra (NLA) is a cornerstone of applied mathematics, scientific computing, and data science. Throughout its long history, NLA algorithmic development has been driven by two competing imperatives, accuracy and efficiency, with the extant challenge to strike the right balance between them.
Randomization has always played a role in NLA, for instance in sometimes providing a good starting point for a Krylov iteration. Another interesting early result was a proposal to randomly precondition linear systems to avoid the need for pivoting, which increases practical speed by reducing data movement parker1995random. However, randomized methods formed a small niche within the mainstream literature; one may speculate that concerns about reproducibility played a role, and perhaps also concerns that randomized methods could not be executed in a high accuracy mode, since that is a characteristic of Monte Carlo methods based on the central limit theorem. (For additional discussion of the role of randomization in NLA, we refer to Kannan2017; Mahoney2011Randomized; martinsson2020.) This survey is concerned with a set of techniques that we refer to as randomized numerical linear algebra (RNLA) that emerged in the early 2000s, and have in the years since attracted significant interest.
@@
RNLA in the 21st century: By the early 2000s, randomization had led to major developments in ground-breaking applications, such as the notable Google PageRank algorithm BrinPage1998Anatomy; HenzingerHeydonMitzenmacherNajork1999; PageBrinMotwaniWinograd1998. Randomized sampling was also being used to compute low-rank matrix decompositions Achlioptas2001; Frieze2004; papadimitriou2000latent.
@@
Over the past two decades, RNLA has developed into a firmly established research area, garnering widespread attention from both practitioners and theorists. In 2006, a three-part series of papers laid the theoretical groundwork for randomized sampling in approximate matrix multiplication DrineasKannanMahoney2006, low-rank matrix decomposition (QR and SVD) DrineasKannanMahoney2006b, and CUR decomposition DrineasKannanMahoney2006c. Around the same time, DeshpandeRademacherVempalaWang2006; Deshpande2006AdaptiveSA investigated adaptive sampling for low-rank approximation; sarlos2006improved proposed a relative error SVD algorithm that utilized randomized subspace embeddings; and 2006_martinsson_random1_orig introduced what would be later referred to as the “randomized SVD (RSVD)”.
@@
Numerical analysts and theoretical computer scientists also provided compelling empirical evidence of the improved performance of RNLA algorithms for practical applications Deshpande2006AdaptiveSA; Friedland2005FastML; Frieze2004; Liberty2007; martinsson2006interpolation; RokhlinTygert2008. These results inspired subsequent investigations of randomized algorithms expressly for linear algebra tasks, such as preconditioning or rangefinding, paired with strong probabilistic guarantees Avron2010; HalkoMartinssonShkolniskyTygert2011; halko2011; MartinssonRokhlinTygert2011; Rokhlin2010. Twenty years later, the scope of research in RNLA now encompasses nearly every aspect of linear algebra, including eigenvalue problems BravermanKrishnanMusco2022; KressnerPlestenjak2024; Tropp2022; general linear solvers BalabanovGrigori2025; ChenEpperlyTroppWebber2025; NakatsukasaTropp2024; orthogonalization BalabanovGrigori2022; FukayaKannanNakatsukasaYamamotoYanagisawa2020; JangGrigori2025; matrix function approximation AmselChenGreenbaumMuscoMusco2024; CortinovisKressnerNakatsukasa2024; PerssonKressner2023; trace estimation HallmanIpsenSaibaba2023; PerssonCortinovisKressner2022; UbaruSaad2018; and many, many more.
@@
Why the R in RNLA? The rise in popularity of randomized algorithms is due in large part to the commensurate progress made in computing technology. In June 2006, the performance of the world’s top supercomputer BlueGene/L was measured at 2.806 Peta-flops/s (floating point operations per second) on the High Performance Linpack (HPL) benchmark, a program that solves a large dense system of linear equations by LU with partial pivoting. As of November 2025, that rate has grown to 1.742 Exa-flops/s, achieved by today’s top supercomputer El Capitan.
@@
Modern computing architectures have made it possible to solve increasingly large-scale problems; however, as evidenced by both theory and practice, the performance of classical algorithms is limited by asymptotic complexity and by the cost of data movement. As a point of reference, the amount of data collected by the Event Horizon Telescope Collaboration to generate the first image of a black hole in 2019 was over 5 Petabytes, which is almost the full storage capacity of El Capitan. In other words, classical matrix algorithms with cubic cost in the input size, such as orthogonalization, are untenable given the large-scale problems now considered in scientific investigations. For the best practical performance, it is necessary to develop algorithms and implementations that can fully leverage computing resources.
@@
Perhaps the greatest benefit of randomization is that it gives us the ability to restructure algorithms to attenuate computational demands on whichever computing resources most hinder performance, such as flops, matrix access, or data movement. For example, randomization admits algorithms that require just a single pass over the input data LiYin2020SinglePassLU; Tropp2018-MorePracticalSketching; TroppStreaming2019, which is especially advantageous for large-scale problems or streaming data. Randomization can also significantly reduce the number of arithmetic operations required by classical algorithms; e.g., for dense $m\times n$ overdetermined least squares problems with $m\gg n$, randomization yields solutions in roughly $O(mn+n^{3})$ operations vs. $O(mn^{2})$ with classical approaches Epperly2024; LacottePilanci2021; OzaslanPilanciArikan2019; RokhlinTygert2008. In short, randomized algorithms produce accurate solutions with high probability while reducing working storage and communication costs, offering improved asymptotic complexity, admitting straightforward parallelization and GPU acceleration, and enabling practical investigations that would be otherwise impossible. For an expanded discussion, see martinsson2020.
@@
RNLA for tensors: The successful incorporation of randomization into classical matrix algorithms inspired similar work to exploit randomization to accelerate tensor computations. Originally proposed in 1927 as a means of working with multi-way physical and chemical measurements Hitchcock1927TheEO; Hitchcock1927Multiple, tensors are now ubiquitous in scientific computing and data science. Applications that involve tensorial data include 3D image reconstruction Bendory2023Autocorrelation; Ghosh2022; Zhang2024MomentMetrics; signal processing Chen2021_WirelessComm; Miron2020; Sidiropoulos2017; quantum chemistry and physics Bischoff2012ManyBodyWaveFunctions; Peng2020; Pierce2021; machine learning Aidini2023FewShotTensorCompletion; Ji2019_ML; Panagakis2024; high-dimensional partial differential equations (PDEs) BoelensVenturiTartakovsky2018; DektorRodgersVenturi2021; Khoromskij2015TensorPDEs; among many others. For detailed surveys of tensors and their applications, we refer the reader to Ballard_Kolda_2025; FriedlandTammali2015LowRankApprox; Grasedyck2013; KoldaBader2009; papalexakis2017.
@@
Unfortunately, tensor practitioners suffer harsh consequences of the curse of dimensionality. For a $d$-way array of uniform size $n$, the storage costs scale as $O(n^{d})$, and most algorithms for analyzing the data scale significantly worse. The key contribution of randomization in this context has been to reveal new algorithms whose complexity scale linearly or close to linearly with the amount of input data, and that involve minimal data movement. This work was inspired by the foundational works CaiafaCichocki2009; delaVega2005; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR from the late 2000s. Two decades later, tensor practitioners increasingly rely on randomization for efficient tensor computations, backed by the strong accuracy guarantees of well-established randomized matrix algorithms. Randomized algorithms have been developed for nearly every type of low-rank tensor decomposition in the literature, including the CP format MaSolomonik2021FastAccurateRandomizedTensor; VervlietDeLathauwer2016RandomizedBlockSampling; Wang2023, the Tucker format AhmadiAsl2021Randomized; CheWeiYan2025EfficientTucker; KressnerPerisa2017RecompressionHadamard, the Tensor Train format CheWeiYan2024; HuberSchneiderWolf2017RandomizedTT_SVD; LiYuBatselier2022FasterTensorTrainSparse, as well as other specialized representations BallaniGrasedyckKluge2013BlackBoxHTucker; BatselierYuDanielWong2018TensorNetworkRandomizedSVD; CastilloHaddockHartsockEtAl2025RandomizedKaczmarzTPRODUCT; Robeva2016; TarzanaghMichailidis2018; WangCuiLi2024SVD_TensorWheel; Zhou2014NTD.
@@
Scope: Our intent with this survey is to bring attention to recent work extending the ideas of RNLA to the world of tensors. We have chosen to include a thorough review of the matrix case in order to make the survey self-contained, and to introduce notation and ideas that will be brought to bear in the tensor regime. (This also provided us the opportunity to review some recent advances on topics such as “fast” random embeddings.) To keep the survey focused, we had to omit several important and interesting topics. For matrices, we do not consider the vast body of work on randomized algorithms for eigenvalues and eigenvalue problems (see, e.g., HeKressnerPlestenjak2025; NakatsukasaTropp2024; Schneider2024_phd) or optimization (see, e.g., derezinski2020debiasing; na2023hessian). We also suppress the details of most theoretical results in favor of more background material; see, e.g., Camano2025_OSI; Kannan2017; martinsson2020. We refer the reader to Murray2023RandNLA for a software-focused survey of RNLA, and to DerezinskiMahoney2024 for a survey of RNLA from the perspective of machine learning. For tensors, we do not consider the tensor train format or other more specialized tensor decompositions and refer the reader to Ballard_Kolda_2025; Grasedyck2013.
Outline: Sect. 24 focus on the matrix setting, beginning with a review of requisite linear algebra material and classical matrix decomposition algorithms. We next survey existing methods of randomized dimension reduction in addition to promising new developments in randomized sketching and sampling. We end the first part of our manuscript by discussing recent work on randomized algorithms for low-rank matrix decompositions, which are the foundation of our randomized algorithms for tensors. Sect. 5 contains the necessary tensor prerequisites, as well as overviews of the two major tensor formats that we consider in our work: CP and Tucker. We then summarize key developments in randomized algorithms to compute low-rank tensor decompositions in these formats in Sect. 6. We close with a few final thoughts on RNLA in Sect. 7.
@@ 2 Matrix Preliminaries
We begin by summarizing notation and key concepts in linear algebra that are frequently referenced in our work. We introduce our notational conventions in Sect. 2.1 and summarize several fundamental matrix decompositions, as well as algorithms to compute them, in Sect. 2.2-2.5.
@@ 2.1 Notation
To refer to coordinates of vectors, we use parentheses with subscripts, i.e. $(\bm{\mathsf{x}})_{i}$ refers to the $i$th coordinate of $\bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptn}$. We also adopt the notation of Golub and Van Loan golub2013 to reference submatrices. Namely, if $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, and $I=[i_{1},i_{2},\dots,i_{k}]\subseteq[m]$ and $J=[j_{1},j_{2},\dots,j_{l}]\subseteq[n]$ are (row and column, resp.) index sets, then $\bm{\mathsf{A}}(I,J)$ denotes
$\bm{\mathsf{A}}(I,J)=\begin{bmatrix}\bm{\mathsf{A}}(i_{1},j_{1})&\bm{\mathsf{A}}(i_{1},j_{2})&\dots&\bm{\mathsf{A}}(i_{1},j_{l})\\ \vdots&\vdots&&\vdots\\ \bm{\mathsf{A}}(i_{k},j_{1})&\bm{\mathsf{A}}(i_{k},j_{2})&\dots&\bm{\mathsf{A}}(i_{k},j_{l})\\ \end{bmatrix}\in\mathbb{F}^{k\times\ell}.$